a scientist uses a gamma degree scale for measuring temperature . in this scale water melts at 10° and boils at 130°. the scientist measures the temperature at which sodium melts to be 127°. express this in kelvin
1
Expert's answer
2017-12-07T10:22:07-0500
We know that water within 100 K: melts at 273,2 K and boils at 373,2 K. From the task we know water melts and boils within 120°. That’s why we may say:
1K = 120° / 100 K = 1.2° or
1° = 100 K / 120° = 0.83 K
The difference between water boiling point and sodium melting point is 3°. Find this difference in kelvin:
ΔT = (3×100)/120 = 2.5 K
The melting point of sodium is:
Tmelt = 373,2 - 2.5 = 370.7 K